OpenStudy (anonymous):

Which fractions are equivalent to 6/8?

OpenStudy (anonymous):

4/6 i think

OpenStudy (chosenmatt):

what can go into 6/8?

sammixboo (sammixboo):

You can take any whole number and multiply it by the numerator and denominator in the fraction \(\rm \color{#20bd23}{\dfrac{6}{8}}\) and those fractions are the same as \(\rm \color{#20bd23}{\dfrac{6}{8}}\) As long as the number you are multiplying by the denominator and numerator is the same

OpenStudy (chosenmatt):

i don't think it's 4/6

sammixboo (sammixboo):

You can also divide \(\rm \color{#20bd23}{\dfrac{6}{8}}\) by 2, and that is also an equivalent fraction

sammixboo (sammixboo):